OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 12x = -5

OpenStudy (anonymous):

(x+6)^2?

OpenStudy (anonymous):

i need a single number bro

OpenStudy (anonymous):

(x+6)^2 = x^2 +12x +36 Compare it to what you have....

OpenStudy (anonymous):

36? sorry im realllyyyy bad at math

OpenStudy (anonymous):

(x+6)^2 = x^2 +12x +36 U have..... x^2 + 12x + 5

OpenStudy (anonymous):

so what is the answerrrr?